OpenStudy (anonymous):

Compound Y has a solubility of 2.0 g/100 mL in water and 20.0 g/100 mL in ether. What weight of compound Y would be removed from a solution of 1.8 g of Y in 100 mL of water by a single extraction with 100 mL of ether? Given the equation: KD= (C1/V1)/(C2/V2)

OpenStudy (aaronq):

compound Y has 10-fold solubility in ether compared to that in water, \(\large 10=\dfrac{(\frac{x}{100~mL})}{(\frac{1.8-x}{100mL})}\) After 1 extraction, mass of compound Y in ether =1.64 g (i haven't done this in a while, but i think it's right)